# Differentially Private Synthetic Data via APIs 3: Using Simulators Instead of Foundation Models

### v1

@@ Abstract
Differentially private (DP) synthetic data, which closely resembles the original private data while maintaining strong privacy guarantees, has become a key tool for unlocking the value of private data without compromising privacy. Recently, Private Evolution (PE) has emerged as a promising method for generating DP synthetic data. Unlike other training-based approaches, PE only requires access to inference APIs from foundation models, enabling it to harness the power of state-of-the-art models. However, a suitable foundation model for a specific private data domain is not always available. In this paper, we discover that the PE framework is sufficiently general to allow inference APIs beyond foundation models. Specifically, we show that simulators—such as computer graphics-based image synthesis tools—can also serve as effective APIs within the PE framework. This insight greatly expands the applicability of PE, enabling the use of a wide variety of domain-specific simulators for DP data synthesis. We explore the potential of this approach, named Sim-PE, in the context of image synthesis. Across three diverse simulators, Sim-PE performs well, improving the downstream classification accuracy of PE by up to 3$\times$ and reducing the FID score by up to 80%. We also show that simulators and foundation models can be easily leveraged together within the PE framework to achieve further improvements. The code is open-sourced in the Private Evolution Python library: https://github.com/microsoft/DPSDA.
private evolution, synthetic data, simulators, differential privacy
@@ 1 Introduction
Leaking sensitive user information is a significant concern in data-driven applications. A common solution is to generate differentially private (DP) (Dwork et al., 2006) synthetic data that closely resembles the original while ensuring strict privacy guarantees. This DP synthetic data can serve as a substitute for the original in various applications, such as model fine-tuning, statistical analysis, and data sharing, while preserving user privacy.
@@
Private Evolution (PE) (Lin et al., 2023; Xie et al., 2024) has recently emerged as a promising method for generating DP synthetic data. PE starts by probing a foundation model to generate random samples, then iteratively selects those most similar to the private data and uses the model to generate more samples that resemble them. Unlike previous state-of-the-art approaches that require fine-tuning open-source models, PE relies solely on model inference. Therefore, PE can be up to 66$\times$ faster than training-based methods (Xie et al., 2024). More importantly, this enables PE to easily harness the cutting-edge foundation models like GPT-4 (OpenAI, 2023) and Stable Diffusion (Rombach et al., 2022), achieving state-of-the-art performance across multiple image and text benchmarks (Lin et al., 2023; Xie et al., 2024; Hou et al., 2024; Zou et al., 2025).
@@
However, PE relies on foundation models suited to the private data domain, which may not always be available. When the model’s distribution significantly differs from the private data, PE’s performance lags far behind training-based methods (Gong et al., 2025).
@@
To address this question, we note that in the traditional synthetic data field—where private data is not involved—domain-specific, non-neural-network simulators remain widely used, especially in domains where foundation models struggle. Examples include computer graphics-based simulators for images, videos, and 3D data (e.g., Blender (Community, 2018) and Unreal (Epic Games, )), physics-based simulators for robotics data (e.g., Genesis (Authors, 2024)), and network simulators for networking data (e.g., ns-2 (Issariyakul et al., 2009)). While these simulators have been successful, their applications in DP data synthesis remain underexplored. This is understandable, as adapting these simulators to fit private data in a DP fashion requires non-trivial, case-by-case modifications. Our key insight is that PE only requires two APIs: RANDOM_API that generates random samples and VARIATION_API that generates samples similar to the given one. These APIs do not have to come from foundation models! Thus, we ask: Can PE use simulators in place of foundation models? If viable, this approach could greatly expand PE’s capabilities and unlock the potential of a wide range of domain-specific simulators for DP data synthesis.
@@
In this paper, we explore this potential in the context of images. We consider two types of simulator access: (1) The simulator is accessible. In this case, we define RANDOM_API as using random simulator parameters to render an image, and VARIATION_API as slightly perturbing the simulator parameters of the given image. (2) The simulator is not accessible—only its generated dataset is released. This scenario is quite common (Wood et al., 2021; Bae et al., 2023), especially when simulator assets are proprietary (Kar et al., 2019; Devaranjan et al., 2020). In this case, we define RANDOM_API as randomly selecting an image from the dataset, and VARIATION_API as randomly selecting a nearest neighbor of the given image. We demonstrate that the resulting algorithm, Sim-PE, outperforms PE with foundation models. Our key contributions are:
@@
    Insight: We identify that PE is not limited to foundation models, making it the first framework capable of utilizing both state-of-the-art foundation models and simulators for DP data synthesis.
@@
    Algorithm: We propose Sim-PE, an extension of PE using simulators, applicable in both scenarios where the simulator or the generated dataset is available. Additionally, we introduce the use of both foundation models and simulators interchangeably during the data synthesis process, allowing for the benefits of both to be leveraged through PE’s easy and standardized interface.
@@
    Results: We demonstrate promising results with Sim-PE. For instance, on the MNIST dataset with $\epsilon=1$, downstream classification accuracy increases to 89.1%, compared to 27.9% with the original PE. Furthermore, combining foundation models with weak simulators results in improved performance compared to using either one alone.
@@ 2.1 Preliminaries
Synthetic data refers to “fake” data generated by models or software for various applications, including data augmentation, model training, and software testing (Lin, 2022). One approach involves machine learning models, ranging from simple statistical models like Gaussian mixtures to more advanced deep neural network-based generative models such as GANs (Goodfellow et al., 2020), diffusion models (Sohl-Dickstein et al., 2015), and auto-regressive models (OpenAI, 2023; Liu et al., 2024a). The other approach relies on simulators. In this paper, we broadly define simulators as non-neural-network data synthesizers with hard-coded, interpretable logic. For example, given network configurations, ns-2 (Issariyakul et al., 2009) can simulate a network and generate network packets. Similarly, given 3D models and lighting configurations, Blender (Community, 2018) can render images and videos of objects. These simulators are widely used across various domains, particularly in cases where the underlying data distribution is too complex for machine learning models to learn effectively.
@@
DP synthetic data requires the synthetic data to be close to a given private dataset, while having a strict privacy guarantee called Differential Privacy (DP) (Dwork et al., 2006). Formally, a mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for any two neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ (i.e., $\mathcal{D}^{\prime}$ has one extra entry compared to $\mathcal{D}$ or vice versa) and for any set $S$ of outputs of $\mathcal{M}$, we have $\mathbb{P}\left(\mathcal{M}\left(\mathcal{D}\right)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{M}\left(\mathcal{D}^{\prime}\right)\in S\right)+\delta$. Smaller $\epsilon$ and $\delta$ imply stronger privacy guarantees. Current state-of-the-art DP image and text synthesis methods rely on machine learning models and typically require model training (Lin et al., 2020a; Beaulieu-Jones et al., 2019; Dockhorn et al., 2022; Yin et al., 2022; Yu et al., 2021; He et al., 2022; Li et al., 2021; Ghalebikesabi et al., 2023a; Yue et al., 2022; Jordon et al., 2019; Harder et al., 2023, 2021b; Vinaroz et al., 2022; Cao et al., 2021).
Private Evolution (PE) (Lin et al., 2023; Xie et al., 2024) is a recent training-free framework for DP data synthesis. PE only requires inference access to the foundation models. Therefore, unlike prior training-based methods, PE can leverage the state-of-the-art models even if they are behind APIs (e.g., GPT-4) and is more computationally efficient.
@@
In more detail, PE has achieved state-of-the-art performance on several image and text benchmarks (Lin et al., 2023; Xie et al., 2024). When using similar open-source pre-trained models, PE attains an image quality score of FID $\leq 7.9$ on CIFAR10 with a privacy cost of $\epsilon=0.67$, a significant improvement over the previous state-of-the-art, which required $\epsilon=32$ (Lin et al., 2023). Furthermore, PE can be up to 66$\times$ more efficient than training-based methods on DP text generation (Xie et al., 2024). By leveraging state-of-the-art models behind APIs—where training-based methods are not applicable—PE further enhances performance, outperforming all prior approaches in downstream text classification accuracy on the OpenReview dataset (Xie et al., 2024). Additionally, PE can be applied in federated learning to shift model training from devices to central servers in a differentially private and more efficient manner (Hou et al., 2024; Zou et al., 2025).
PE is versatile across data modalities, as long as suitable foundation models are available with two functions: (1) RANDOM_API that generates a random sample (e.g., a random bird image), and (2) VARIATION_API that generates slight modifications of the given sample (e.g., a similar bird image). PE works by first calling RANDOM_API to get an initial set of synthetic samples, and then iteratively refine this set by selecting the closest ones to the private samples (in a DP manner) and calling VARIATION_API to generate more of such samples. The full PE algorithm from Lin et al. (2023) is attached in App. A for completeness.
@@ 2.2 Motivation
While PE achieves state-of-the-art performance on several image and text benchmarks (Lin et al., 2023; Xie et al., 2024; Hou et al., 2024), its performance significantly drops when there is a large distribution shift between the private data and the foundation model’s pre-trained data (Gong et al., 2025). For instance, when using the MNIST dataset (LeCun, 1998) (handwritten digits) as the private data, training a downstream digit classifier (10 classes) on DP synthetic data (with $\epsilon=1$) from PE—using a foundation model pre-trained on ImageNet—yields an accuracy of only 27.9%. Since relevant foundation models may not always be available for every domain, this limitation hinders PE’s applicability in real-world scenarios. Extending PE to leverage simulators could significantly expand its potential applications.
More broadly, as discussed in § 2.1, simulators cannot be substituted by foundation models in (non-DP) data synthesis across many domains. Unfortunately, current state-of-the-art DP synthetic data methods are deeply reliant on machine learning models (e.g., requiring model training) and cannot be applied to simulators. By extending PE to work with simulators, we aim to unlock the potential of simulators in DP data synthesis.
@@ 3.1 Overview
In this paper, we focus on DP image generation. The beauty of the Private Evolution framework is that it isolates DP mechanism from data generation backend. In particular, any data generation backend that supports RANDOM_API and VARIATION_API can be plugged into the framework and transformed into a DP data synthesis algorithm. Therefore, our goal is to design RANDOM_API and VARIATION_API for image simulators.
We notice that existing popular image simulators provide different levels of access. Some simulators are open-sourced. Examples include Kubric (Greff et al., 2022), a Blender-based renderer for multi-object images/videos; 3D Teapot (Lin et al., 2020b; Eastwood & Williams, 2018), an OpenDR-based renderer for teapot images; and python-avatar (Escartín, 2021), a rule-based generator for avatars. However, the assets (e.g., 3D models) used in these renderers are often proprietary. Therefore, many simulator works choose to release only the generated datasets without the simulator code. Examples include the Face Synthetics (Wood et al., 2021) and the DigiFace-1M (Bae et al., 2023) datasets, both generated using Blender-based renderers for human faces. In § 3.2 and 3.3, we discuss the design for simulators with code access and data access, respectively.
Moreover, since simulators and foundation models provide the same RANDOM_API and VARIATION_API interfaces to PE, we can easily utilize both together in the data generation process. § 3.4 discusses the methedology.
Privacy analysis. Since we only modify RANDOM_API and VARIATION_API, the privacy guarantee is exactly the same as Lin et al. (2023), and we skip it here.
@@ 3.2 Sim-PE with Simulator Access
While different simulators have very different programming interfaces, most of them can be abstracted in the same way. Given a set of $p$ categorical parameters $\xi_{1},\ldots,\xi_{p}$ and $q$ numerical parameters $\phi_{1},\ldots,\phi_{q}$ where $\xi_{i}\in\Xi_{i}$ and $\phi_{i}\in\Phi_{i}$, the simulator $\mathcal{S}$ generates an image $\mathcal{S}\left(\xi_{1},\ldots,\xi_{p},\phi_{1},\ldots,\phi_{q}\right)$. For example, for face image renders (Wood et al., 2021; Bae et al., 2023), $\xi_{i}$s could be the ID of the 3D human face model and the ID of the hair style, and $\phi_{i}$s could be the angle of the face and the strength of lighting.
For RANDOM_API, we simply draw each parameter randomly from its corresponding feasible set. Specifically, we define
@@
$\quad\quad=\mathcal{S}\left(\xi_{1}^{\prime},\ldots,\xi_{p}^{\prime},\phi_{1}^{\prime},\ldots,\phi_{q}^{\prime}\right),$ (2)
$\text{where }\phi_{i}^{\prime}\sim\text{Uniform}\left(\left[\phi_{i}-\alpha,\phi_{i}+\alpha\right]\cap\Phi_{i}\right)\text{ and }$
$\quad\xi_{i}^{\prime}=\begin{cases}\text{Uniform}\left(\Xi_{i}\right),&\text{with probability }\beta\\ \xi_{i},&\text{with probability }1-\beta\end{cases}.$

### v2

@@ Abstract
Differentially private (DP) synthetic data, which closely resembles the original private data while maintaining strong privacy guarantees, has become a key tool for unlocking the value of private data without compromising privacy. Recently, Private Evolution (PE) has emerged as a promising method for generating DP synthetic data. Unlike other training-based approaches, PE only requires access to inference APIs from foundation models, enabling it to harness the power of state-of-the-art (SoTA) models. However, a suitable foundation model for a specific private data domain is not always available. In this paper, we discover that the PE framework is sufficiently general to allow APIs beyond foundation models. In particular, we demonstrate that many SoTA data synthesizers that do not rely on neural networks—such as computer graphics-based image generators, which we refer to as simulators—can be effectively integrated into PE. This insight significantly broadens PE’s applicability and unlocks the potential of powerful simulators for DP data synthesis. We explore this approach, named Sim-PE, in the context of image synthesis. Across four diverse simulators, Sim-PE performs well, improving the downstream classification accuracy of PE by up to 3$\times$, reducing FID by up to 80%, and offering much greater efficiency. We also show that simulators and foundation models can be easily leveraged together within PE to achieve further improvements. The code is open-sourced in the Private Evolution Python library: https://github.com/microsoft/DPSDA.
Main updates in arXiv V2: Major revision with additional experiments and improved writing.
@@ 1 Introduction
Leaking sensitive user information is a major concern in data-driven applications. A common solution is to generate differentially private (DP) [15] synthetic data that resembles the original while ensuring strong privacy guarantees. Such data can substitute the original in tasks like model fine-tuning, statistical analysis, and data sharing, while preserving user privacy [7, 43, 58, 34].
@@
Private Evolution (PE) [46, 62] has recently emerged as a promising method for DP data synthesis. It begins by probing a foundation model to produce random samples, then iteratively selects those most similar to private data and uses the model to generate more like them. Unlike prior state-of-the-art (SoTA) methods that fine-tune open-source models, PE relies solely on model inference–making it up to 66$\times$ faster [62]. More importantly, this allows PE to easily leverage cutting-edge foundation models like GPT-4 [52] and Stable Diffusion [54], achieving SoTA performance on multiple image and text benchmarks [46, 62, 32, 71, 33]. PE has also been adopted in industry [2, 1].
@@
However, PE relies on foundation models suited to the private data domain, which may not always be available. When the model’s distribution significantly differs from the private data, PE’s performance lags far behind training-based methods [22].
@@
To address this question, we note that in the traditional synthetic data field—where private data is not involved—non-neural-network data synthesizers remain widely used, especially in domains where foundation models struggle. Examples include computer graphics-based renders for images, videos, and 3D data (e.g., Blender [11] and Unreal [18]), physics-based simulators for robotics data (e.g., Genesis [3]), and network simulators for networking data (e.g., ns [35, 53]). For brevity, we refer to these tools as simulators. While these simulators have been successful, their applications in DP data synthesis remain underexplored. This is understandable, as adapting these simulators to fit private data in a DP fashion requires non-trivial, case-by-case modifications. Our key insight is that PE only requires two APIs: RANDOM_API that generates random samples and VARIATION_API that generates samples similar to the given one. These APIs do not have to come from foundation models! Thus, we ask: Can PE use simulators in place of foundation models? If viable, this approach could greatly expand PE’s capabilities and unlock the potential of a wide range of simulators for DP data synthesis.
Figure 1: Unlike prior PE work that relies solely on foundation models, we show that PE is also compatible with non-neural-network data synthesis tools, which we call simulators. This greatly broadens PE’s applicability and enables SoTA simulators for DP data synthesis.
@@
In this paper, we propose Sim-PE (Fig. 1) to exploit this potential for image generation. We consider two types of simulator access: (1) The simulator is accessible. We define RANDOM_API as rendering an image with random simulator parameters, and VARIATION_API as slightly perturbing the simulator parameters of the given image. (2) The simulator is inaccessible—only its generated data is released. This scenario is quite common [61, 4], especially when simulator assets are proprietary [38, 12]. In this case, we define RANDOM_API as randomly selecting an image from the dataset, and VARIATION_API as randomly selecting a nearest neighbor of the given image. We demonstrate that with suitable simulators, Sim-PE can outperform PE with foundation models. Our key contributions are:
@@
    Advancing PE. We discover that PE can leverage tools beyond foundation models and propose Sim-PE—an extension that uses simulators, significantly broadening PE’s applicability. We also introduce the use of both foundation models and simulators interchangeably during synthesis, allowing for the benefits of both to be leveraged through PE’s easy and standardized interface.
@@
    Bringing simulators to DP synthetic data. Although simulators are widely used and powerful (App. A), they have been largely absent from DP data generation. Sim-PE is the first framework to unlock their potential in this space.
@@
    Results. We demonstrate promising results with Sim-PE. For instance, on the MNIST dataset with $\epsilon=1$, downstream classification accuracy increases to 89.1%, compared to 27.9% with the original PE. Furthermore, combining foundation models with weak simulators results in improved performance compared to using either one alone.
@@ 2.1 Preliminaries
Synthetic data refers to “fake” data generated by models or software for various applications, including data augmentation, model training, and software testing [43]. While neural-network-based generative models such as GANs [23], diffusion models [56], and auto-regressive models [52, 49] are widely used, non-neural-network tools remain SoTA in many applications. For example, ns [35, 53] can simulate networks and generate network packets based on network configurations. Blender [11], given 3D models and lighting configurations, can render images and videos, and is extensively used in movie production. In this paper, we refer to these tools as simulators. See App. A for a discussion on the continued importance of simulators, even in the era of large foundation models.
@@
DP synthetic data requires the synthetic data to be close to a given private dataset, while having a strict Differential Privacy (DP) [15] guarantee. Formally, a mechanism $\mathcal{M}$ is $(\epsilon,\delta)$-DP if for any two neighboring datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ (i.e., $\mathcal{D}^{\prime}$ has one extra entry compared to $\mathcal{D}$ or vice versa) and for any set $S$ of outputs of $\mathcal{M}$, we have $\mathbb{P}\left(\mathcal{M}\left(\mathcal{D}\right)\in S\right)\leq e^{\epsilon}\mathbb{P}\left(\mathcal{M}\left(\mathcal{D}^{\prime}\right)\in S\right)+\delta$. Smaller $\epsilon$ and $\delta$ imply stronger privacy guarantees. Current SoTA DP image and text synthesis methods typically requires neural network training [47, 5, 14, 66, 68, 29, 42, 21, 69, 37, 28, 27, 60, 8, 10].
@@
Private Evolution (PE) [46, 62] is a recent training-free framework for DP data synthesis. PE only requires inference access to the foundation models. Therefore, unlike prior training-based methods, PE can leverage the SoTA models even if they are behind APIs (e.g., GPT-4) and is more computationally efficient [46, 62]. In more detail, PE has achieved SoTA performance on several image and text benchmarks [46, 62]. When using similar open-source pre-trained models, PE attains an image quality score of FID $\leq 7.9$ on CIFAR10 with a privacy cost of $\epsilon=0.67$, a significant improvement over the previous SoTA, which required $\epsilon=32$ [46]. Furthermore, PE can be up to 66$\times$ more efficient than training-based methods on DP text generation [62]. By leveraging SoTA models behind APIs—where training-based methods are not applicable—PE further enhances performance, outperforming all prior approaches in downstream text classification accuracy on the OpenReview dataset [62]. Additionally, PE can be applied in federated learning to shift model training from devices to central servers in a differentially private and more efficient manner [32, 71, 33]. Moreover, PE has been adopted by some of the largest IT companies [2, 1].
@@ 2.2 Motivation
While PE achieves SoTA performance on several image and text benchmarks [46, 62, 32], its performance significantly drops when there is a large distribution shift between the private data and the foundation model’s pre-trained data [22]. For instance, when using the MNIST dataset [40] (handwritten digits) as the private data, training a downstream digit classifier (10 classes) on DP synthetic data (with $\epsilon=1$) from PE—using a foundation model pre-trained on ImageNet—yields an accuracy of only 27.9%. Since relevant foundation models may not always be available for every domain, this limitation hinders PE’s applicability in real-world scenarios. Extending PE to leverage simulators could significantly expand its potential applications.
More broadly, as discussed in § 2.1 and A, simulators cannot be substituted by foundation models in (non-DP) data synthesis across many domains. Unfortunately, current SoTA DP synthetic data methods are deeply reliant on machine learning models (e.g., requiring model training) and cannot be applied to simulators. By extending PE to work with simulators, we aim to unlock the potential of simulators in DP data synthesis.
@@ 3 Sim-PE: Private Evolution (PE) with Simulators
In this paper, we focus on DP image generation. A key advantage of the PE framework is that it decouples the DP mechanism from the data generation backend. Specifically, any backend that supports (1) RANDOM_API, which generates a random sample (e.g., a random bird image), and (2) VARIATION_API, which produces slight variations of a given sample (e.g., a similar bird image), can be integrated into PE and turned into a DP data synthesis algorithm. Prior work on PE [46, 62, 32, 71, 33, 57] has exclusively used foundation models to implement these APIs. Our key insight is that these APIs do not need to be powered by foundation models: traditional data synthesizers that do not rely on neural networks—referred to as simulators—can also be used to implement RANDOM_API and VARIATION_API.
In the following sections, we first provide an overview of the Sim-PE algorithm (§ 3.1) and the design of its APIs (§ 3.2 and 3.3), then discuss how simulators and foundation models can be jointly used in PE to leverage the strengths of both (§ 3.4).
@@ 3.1 Overview
Algorithm overview. Except for the APIs, Sim-PE largely follows the same workflow as PE. For completeness, we briefly describe the workflow here and include the full algorithm in App. B. We first use RANDOM_API to generate an initial set of random samples (Alg. 1). Then, we iteratively refine these samples using the private data. In each iteration:
  • $\bullet$
    Each private sample casts a vote for its closest synthetic sample. This yields a histogram (denoted DP_NN_HISTOGRAM) reflecting how well each synthetic sample aligns with the private data (Alg. 1). To ensure differential privacy, Gaussian noise is added to this histogram.
  • $\bullet$
    We sample synthetic data according to the noisy histogram, giving higher likelihood to those samples that align more closely with private data (Algs. 1 and 1).
  • $\bullet$
    We apply VARIATION_API to the drawn samples to generate additional variants (Alg. 1). These samples become the initialization of the next iteration.
The synthetic samples at the final iteration constitute the DP synthetic dataset.
Theoretical analysis. Since we only modify RANDOM_API and VARIATION_API, the privacy guarantee and convergence analysis are exactly the same as PE [46] (more details in App. D).
Considered simulators. Existing popular image simulators often provide different levels of access. Some simulators are open-sourced. Examples include Kubric [25], a Blender-based renderer for multi-object images/videos; 3D Teapot [48, 17], an OpenDR-based renderer for teapot images; and python-avatar [19], a rule-based generator for avatars. However, the assets (e.g., 3D models) used in these renderers are often proprietary. Therefore, many simulator works choose to release only the generated datasets without the simulator code. Examples include the Face Synthetics [61] and the DigiFace-1M [4] datasets, both generated using Blender-based renderers for human faces. In § 3.2 and 3.3, we discuss the design for simulators with code access and data access, respectively.
@@ 3.2 Sim-PE with Simulator Access
While different simulators have very different programming interfaces, most of them can be abstracted in the same way. Given a set of $p$ categorical parameters $\xi_{1},\ldots,\xi_{p}$ and $q$ numerical parameters $\phi_{1},\ldots,\phi_{q}$ where $\xi_{i}\in\Xi_{i}$ and $\phi_{i}\in\Phi_{i}$, the simulator $\mathcal{S}$ generates an image $\mathcal{S}\left(\xi_{1},\ldots,\xi_{p},\phi_{1},\ldots,\phi_{q}\right)$. For example, for face image renders [61, 4], $\xi_{i}$s could be the ID of the 3D human face model and the ID of the hair style, and $\phi_{i}$s could be the angle of the face and the strength of lighting. [^1] For well-documented simulators, obtaining the list of parameters is straightforward. For example, python-avatar, used in § 4.2, lists its parameters in the README. Alternatively, one can use the approach in § 3.3, which does not require explicit parameter identification. Note that numerical parameters may be continuous (e.g., $\in[0,2]$) or discrete (e.g., $\in\left\{0,1,2\right\}$).
For RANDOM_API, we simply draw each parameter randomly from its corresponding feasible set:
@@
$\text{where }\phi_{i}^{\prime}\sim\text{Uniform}\left(\left[\phi_{i}-\alpha,\phi_{i}+\alpha\right]\cap\Phi_{i}\right)\text{ and }\xi_{i}^{\prime}=\begin{cases}\text{Uniform}\left(\Xi_{i}\right),&\text{with probability }\beta\\ \xi_{i},&\text{with probability }1-\beta\end{cases}.$